Session: Sep 1997 (session 3)
Frequency: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (k. Borkowski at Tue Sep 23 17:38:13 METDST 1997) Minor failures
Two scans almost entirely lost: at 9:35 and 10:14, due to problems with the telescope control.

Medicina (Fredrik T Rantakyro at Tue Sep 23 17:51:02 METDST 1997) Success

Westerbork (tony foley at Tue Sep 23 17:58:52 METDST 1997) Minor failures
Read headstack was sometimes stuck Some nonsense commands (sent from barcode reader by accident) about 10:20UT.

Onsala (Antonis Polatidis at Tue Sep 23 20:45:13 METDST 1997) Success
No known problems

Cambridge (Alastair Gunn at Tue Sep 23 22:05:18 METDST 1997) Success
No known problems.

Effelsberg (dgraham at Wed Sep 24 13:27:08 METDST 1997) Success

Noto (C.Trigilio at Thu Sep 25 09:23:11 METDST 1997) Minor failures
Lost the first part of experiment for problems in the communication with the telescope. The problem was resolved only at 7:40 UT. Second part OK.

Calibration (Fredrik T Rantakyro at Wed Nov 5 16:02:40 MET 1997) Minor failures
The ANTAB formatted TSYS information can be found on VLBEER in the SEP97 directory. The file is called v085i.antab and contains the latest antenna calibration information.